OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2). A. 1/5 B. 5 C. -1/5 D. -5

OpenStudy (anonymous):

slope of a line, m = (y1-y2)/(x1-x2) here, m = (8- - 2) / (4 - 6) = 10/(-2) = -1/5

OpenStudy (anonymous):

 (-2-8)/(6-4) = -10/2 = -5

OpenStudy (anonymous):

okay...